# Kandinsky Conformal Prediction:

### v1

@@ 1 Introduction
Conformal prediction (Vovk et al., 2005; Lei and Wasserman, 2014; Papadopoulos et al., 2002b) is a framework for constructing prediction sets with formal coverage guarantees. Given a predictor $f:\cal X\to\cal Y$ and samples from a distribution $\cal D$, the goal is to build a prediction set function ${\cal C}$ that, for a given covariate vector $X$, outputs a subset of $\cal Y$ guaranteed to include the true label $Y$ with a target probability. Ideally, the size of ${\cal C}(X)$ reflects the uncertainty in $f(X)$.
Specifically, using a trained predictor $f$, a calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$, and the covariate vector of a test example $X_{n+1}$, a common objective is to construct ${\cal C}(X_{n+1})$ such that with high probability over the draws of the calibration dataset and any internal randomness of ${\cal C}$, the following marginal coverage guarantee holds:
$\mathbb{P}_{(X_{n+1},Y_{n+1})\sim\cal D}\left[Y_{n+1}\in{\cal C}(X_{n+1})\right]=1-\alpha,$ (1)
for some $\alpha$ in $(0,1)$.111 In Vovk (2012) this type of coverage is called training conditional validity for conformal prediction. Another type of guarantee that is common in the conformal prediction literature is $\mathbb{\mathbb{P}}\left[Y_{n+1}\in{\cal C}(X_{n+1})\right]=1-\alpha$, where the probability is taken over all $n+1$ points and the internal randomness of ${\cal C}$ (e.g. see Gibbs et al. (2023)). This guarantee ensures coverage “on average” over test samples $(X_{n+1},Y_{n+1})$ drawn from $\cal D$. However, a limitation of marginal coverage is that the prediction sets might undercover certain values of $(X_{n+1},Y_{n+1})$ while overcovering others. A potential solution is to require $(1-\alpha)$-coverage conditioning on the value of $(X_{n+1},Y_{n+1})$. Unfortunately, we generally cannot obtain non-trivial prediction sets that satisfy such pointwise coverage (Vovk, 2012; Lei and Wasserman, 2014; Barber et al., 2021).
@@
In this paper, we expand the scope of conditional coverage guarantees by achieving the best of both worlds from these two lines of prior work. Building on Gibbs et al. (2023); Jung et al. (2023), we provide coverage guarantees for overlapping groups while, in the spirit of Mondrian conformal prediction (Vovk et al., 2003), allowing grouping functions to depend jointly on both the covariates $X_{n+1}$ and the label $Y_{n+1}$. More generally, our guarantees hold for fractional groups, where membership is defined as a probabilistic function over $X_{n+1}$ and $Y_{n+1}$. This added flexibility enables the modeling of subgroups based on protected attributes that are not explicitly observed but can be inferred through covariates and labels (Romano et al., 2020a). We name our method Kandinsky Conformal Prediction, drawing inspiration from Wassily Kandinsky’s compositions of overlapping geometric forms (see Figure 2), as we provide conditional coverage guarantees for groups that are flexible, overlapping, and probabilistic.
@@
To complement our main algorithm, which ensures training-conditional validity, we also provide a test-time inference algorithm that obtains expected coverage that also takes randomness over the draws on calibration data. By extending Gibbs et al. (2023)’s approach to accommodate group functions defined over both covariates and labels, we obtain an expected coverage error bound of $\mathcal{O}(n_{{\cal G}}^{-1})$. However, this result comes at the cost of computational efficiency, as the algorithm needs to perform test-time inference that solves an instance of quantile regression for every new test example.
@@
    We evaluate our algorithm on real-world tasks, including income prediction and toxic comment detection. Our results show that Kandinsky Conformal Prediction consistently achieves the best-calibrated conditional coverage while scaling effectively with the number of groups.
@@ 1.1 Related Work
We introduce Kandinsky Conformal Prediction as an extension of the Mondrian approach, providing coverage guarantees for overlapping group structures. Prior work on overlapping group-conditional coverage primarily considers groups defined solely by covariates $X_{n+1}$. Barber et al. (2021) addresses this by assigning the most conservative prediction set to points in multiple groups. Our algorithm builds on the quantile regression techniques of Jung et al. (2023) and Gibbs et al. (2023), which compute per-example thresholds on non-conformity scores. Additionally, Areces et al. (2024) establishes a lower bound on group-conditional coverage error rates. Since we generalize to group functions that depend on both covariates and labels, their bound applies to our setting and shows that our error rate is minimax-optimal.
@@
Finally, our work advances the study of conformal prediction under distribution shifts between calibration and test data. Existing research largely focuses on specific cases, such as covariate shift (Tibshirani et al., 2019; Qiu et al., 2023; Yang et al., 2024; Pournaderi and Xiang, 2024) or label shift (Podkopaev and Ramdas, 2021), where they consider a single target distribution. By introducing a more general class of group functions that depend jointly on covariates and labels, we establish coverage guarantees that remain valid under broader distribution shifts over $\cal X\times\cal Y$, including multiple potential target distributions.
@@ 2 Preliminaries
We consider prediction tasks over ${\cal X}\times{\cal Y}$, where ${\cal X}$ represents the covariate domain and ${\cal Y}$ represents the label domain. ${\cal Y}$ can be either finite for classification tasks or infinite for regression tasks. Unless stated otherwise, we assume that the observed data is drawn from a distribution ${\cal D}$, defined over ${\cal X}\times{\cal Y}$. We use weight functions $w:\cal X\times\cal Y\to\mathbb{R}$ to model our conditional coverage guarantees. Let $(1-\alpha)$ be the coverage level we aim for, where $\alpha\in(0,1)$. For any weight function $w$, we measure the error of a prediction set function ${\cal C}$ by its weighted coverage deviation:
$\operatorname{wCD}({\cal C},\alpha,w)\coloneqq\mathbb{E}[w(X,Y)(\mathbbm{1}\{Y\in{\cal C}(X)\}-(1-\alpha))],$
where the expectation is taken over the test point $(X,Y)$, drawn from ${\cal D}$, and internal randomness of ${\cal C}$. The bounds of $\operatorname{wCD}({\cal C},\alpha,w)$, for a weight function $w$, in our results depend on the quantity $\|w\|_{\infty}\coloneqq\sup_{x\in{\cal X},y\in{\cal Y}}|w(x,y)|$.
@@
Let $\alpha$ be a parameter in $(0,1)$, and let $\{s_{i}\}_{i\in[n]}$ be a set of $n$ points, where for all $i\in[n]$, $s_{i}\in\mathbb{R}$. Then, the largest $\theta^{*}\in\mathbb{R}$ such that
@@ 3 Kandinsky Conformal Prediction
Given the predictor $f$, we use a non-conformity score function $S:{\cal X}\times{\cal Y}\to\mathbb{R}$ to measure how close the prediction $f(x)$ is to the label $y$ for any data point $(x,y)\in{\cal X}\times{\cal Y}$. As is common in conformal prediction, our method ensures the desirable coverage for any score function $S$.
Our method consists of two steps: one is performed during training, and the other at test time. Similar to Jung et al. (2023), the first step is a quantile regression on the scores of the $n$ data points in the calibration dataset, $\{S(X_{i},Y_{i})\}_{i\in[n]}$. However, in our more general framework, we allow for the use of a randomized score function to break potential ties in quantile regression, while keeping the assumptions about distribution ${\cal D}$ minimal. We implement this through a randomized score function ${\widetilde{S}}:{\cal X}\times{\cal Y}\times\mathbb{R}$ that takes as input the covariates, labels and some random noise $\varepsilon\in\mathbb{R}$. Therefore, we work with the set of modified scores $\{{\widetilde{S}}(X_{i},Y_{i},\varepsilon_{i})\}_{i\in[n]}$, where for every point $i$, $\varepsilon_{i}$ is drawn independently from a distribution ${\cal D}_{\text{rn}}$. In Algorithm 1, we compute a $(1-\alpha)$-“quantile function” $\widehat{q}$. This $\widehat{q}$ is a weight function from ${\cal W}$ that minimizes the average pinball loss of the $n$ (randomized) scores with parameter $\alpha$.
@@
${\cal C}(X_{n+1})=\{y:{\widetilde{S}}(X_{n+1},y,\varepsilon_{n+1})\leq\widehat{q}(X_{n+1},y)\}.$
Return: ${\cal C}(X_{n+1})$
Our main result is stated in Theorem 3.1, where we prove that, under the assumption that the distribution of the randomized score ${\widetilde{S}}$ conditioned on the value of the basis $\Phi$ is continuous, the weighted coverage deviation of ${\cal C}$ converges to zero with high probability at a rate of $O(\sqrt{d/n}+d/n)$, where $d$ is the dimension of $\Phi$. When the distribution of $S(X,Y)\mid\Phi(X,Y)$ is continuous, this theorem holds for the original score function $S$, since setting ${\widetilde{S}}(x,y,\varepsilon)=S(x,y)$ satisfies the assumptions.
@@
Let parameters $\alpha,\delta\in(0,1)$, and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n+1]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. There exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$ given by Algorithms 1 and 2 satisfies, for every $w\in{\cal W}$,
$\left|\operatorname{wCD}({\cal C},\alpha,w)\right|\leq\|w\|_{\infty}\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right).$
@@
Let $\alpha,\delta,{\cal W}$ be specified as in Theorem 3.1 and consider the same assumptions on $\{(X_{i},Y_{i})\}_{i\in[n+1]}$, $\widetilde{S}$, and $\Phi$ in Theorem 3.1. Then, there exists an absolute constant $C$ such that the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$\mathbb{E}_{D,E}\left[\sup_{w\in{\cal W}}\frac{\operatorname{wCD}({\cal C},\alpha,w)}{\|w\|_{\infty}}\right]\leq C\sqrt{\frac{d}{n}}+\frac{d}{n},$
@@ 3.1 Applications
Suppose we want to achieve the target coverage $(1-\alpha)$ for every group of points $G\subset{\cal X}\times{\cal Y}$ within a finite set of potentially overlapping groups ${\cal G}$. Let $\pi_{G}({\cal C})$ be the probability that the true label is included in the prediction set ${\cal C}$, conditioned on the datapoint $(X_{n+1},Y_{n+1})$ belonging to some group $G$. Formally, we use the notation
$\pi_{G}({\cal C})\coloneqq\mathbb{P}[Y_{n+1}\in{\cal C}(X_{n+1})\mid(X_{n+1},Y_{n+1})\in G].$
We want to ensure that for every $G\in{\cal G}$, $\pi_{G}({\cal C})=1-\alpha$. Here we can define $\Phi$ to be a $|{\cal G}|$-dimensional vector that has an entry $\mathbbm{1}\{(x,y)\in G\}$ for every $G\in{\cal G}$. Corollary 3.3 provides the rate at which $\pi_{G}({\cal C})$ converges to $(1-\alpha)$, when ${\cal C}$ is constructed by Algorithms 1 and 2. For conciseness, we define $\pi_{G}\coloneqq\mathbb{P}[(X_{n+1},Y_{n+1})\in G]$.
@@
Let parameters $\alpha,\delta\in(0,1)$, and ${\cal W}=\{\sum_{G\in{\cal G}}\beta_{G}\mathbbm{1}\{(x,y)\in G\}:\beta_{G}\in\mathbb{R},\forall G\in{\cal G}\}$. Under the same assumptions on $\{(X_{i},Y_{i})\}_{i\in[n+1]}$, $\widetilde{S}$, and $\Phi$ in Theorem 3.1, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$ given by Algorithms 1 and 2 satisfies, for all $G\in{\cal G}$,
$\left|\mathbb{E}_{\varepsilon_{n+1}}[\pi_{G}({\cal C})]-(1-\alpha)\right|\leq\frac{1}{\pi_{G}}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right).$
@@
A special case of group-conditional conformal prediction is Mondrian conformal prediction, where groups in ${\cal G}$ are disjoint and form a partition of the domain ${\cal X}\times{\cal Y}$. Since each $(X_{i},Y_{i})$ in the calibration dataset belongs to exactly one group in ${\cal G}$, we can verify that Algorithm 1 computes a value $\beta_{G}\in\mathbb{R}$ for every group $G$ independently. By Lemma 2.3, for every $G\in{\cal G}$, if $\beta_{G}$ is the largest value that minimized the pinball loss, then it is a $(1-\alpha)$-quantile of the scores of the calibration datapoints that belong to $G$. As a result, for a test point $X_{n+1}$, we construct the prediction set ${\cal C}(X_{n+1})$ by including every $y$ where the score ${\widetilde{S}}(X_{n+1},y,\varepsilon)$ is at most the quantile $\beta_{G}$ corresponding to the group $G$ that contains $(X_{n+1},y)$. This method is the same as the one presented in Vovk et al. (2003), with the difference that in that method, the quantiles are slightly adjusted to achieve the expected coverage guarantee through exchangeability.
@@
$\pi^{Z}_{G}({\cal C})\coloneqq\mathbb{P}[Y_{n+1}\in{\cal C}(X_{n+1})\mid Z_{n+1}\in G].$
Then, we want to ensure that $\pi^{Z}_{G}({\cal C})=1-\alpha$ for every $G\in{\cal G}$. To achieve this, we set the basis of the weight function class as the probability of $Z\in G$ given a statistic $\phi(X,Y)$. For every $G\in{\cal G}$,
@@
For our theoretical result, we assume that we already know these conditional probabilities, $\Phi_{G}(x,y)$, for all $G\in{\cal G}$, and that the calibration and the test data do not include unobserved attributes. Corollary 3.4 shows that by constructing ${\cal C}$ according to Algorithms 1 and 2, the coverage converges to $(1-\alpha)$ at an $O(\sqrt{|{\cal G}|/n}+|{\cal G}|/n)$ rate. For brevity, we define $\pi_{G}^{Z}\coloneqq\mathbb{P}[Z\in G]$.
@@
Let $\alpha,\delta\in(0,1)$, $\phi(X,Y)$ be a sufficient statistic for ${\widetilde{S}}$, such that ${\widetilde{S}}$ is conditionally independent of $X,Y$ given $\phi$, and ${\cal W}=\{\sum_{G\in{\cal G}}\beta_{G}\Phi_{G}:\beta_{G}\in\mathbb{R},\forall~{}G\in{\cal G}\}.$ Under the same assumptions on $\{(X_{i},Y_{i})\}_{i\in[n+1]}$, $\widetilde{S}$, and $\Phi$ in Theorem 3.1, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$ given by Algorithms 1 and 2 satisfies, for all $G\in{\cal G}$,
$\left|\mathbb{E}_{\varepsilon_{n+1}}[\pi^{Z}_{G}({\cal C})]-(1-\alpha)\right|\leq\frac{1}{\pi_{G}^{Z}}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4/\delta)}{n}}\right).$

### v2

@@ 1 Introduction
Conformal prediction (Vovk et al., 2005; Lei and Wasserman, 2014; Papadopoulos et al., 2002b) is a framework for constructing prediction sets with formal coverage guarantees. Given a predictor $f:\cal X\to\cal Y$ and $n$ samples $\{(X_{i},Y_{i})\}_{i\in[n]}$ from a distribution $\cal D$, the goal is to build a prediction set function ${\cal C}$ that, for a given covariate vector $X_{n+1}$, outputs a subset of $\cal Y$ guaranteed to include the true label $Y$ with a target probability. Ideally, the size of ${\cal C}(X_{n+1})$ reflects the uncertainty in $f(X_{n+1})$.
A common coverage guarantee, termed as training conditional validity Vovk (2012), requires that for any $f$ the training algorithm returns a (potentially randomized) prediction set function ${\cal C}$ such that, with probability $1-\delta$ over the draws of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$, the following holds:
$\mathbb{P}\left[Y_{n+1}\in{\cal C}(X_{n+1})\mid\{(X_{i},Y_{i})\}_{i\in[n]}\right]\approx 1-\alpha,$ (1)
for some $\alpha$ and $\delta$ in (0,1)(0,1)( 0 , 1 ). This guarantee ensures coverage on average over random test samples $(X_{n+1},Y_{n+1})$ drawn from $\cal D$. A weaker guarantee, called the marginal coverage guarantee, requires that $\mathbb{\mathbb{P}}\left[Y_{n+1}\in{\cal C}(X_{n+1})\right]\approx 1-\alpha$, where the probability is taken over all $n+1$ points (Vovk et al., 2005; Gibbs et al., 2023). However, a limitation of such average coverage guarantees is that the prediction sets might undercover certain values of $(X_{n+1},Y_{n+1})$ while overcovering others. A potential solution is to require $(1-\alpha)$-coverage conditioning on the value of $(X_{n+1},Y_{n+1})$. Unfortunately, we generally cannot obtain non-trivial prediction sets that satisfy such pointwise coverage (Vovk, 2012; Lei and Wasserman, 2014; Barber et al., 2021).
@@
In this paper, we expand the scope of conditional coverage guarantees by achieving the best of both worlds from these two lines of prior work. Building on Gibbs et al. (2023); Jung et al. (2023), we provide coverage guarantees for overlapping groups while, in the spirit of Mondrian conformal prediction (Vovk et al., 2003), allowing grouping functions to depend jointly on both the covariates $X_{n+1}$ and the label $Y_{n+1}$. More generally, our guarantees hold for fractional groups, where membership is defined as a probabilistic function over $X_{n+1}$ and $Y_{n+1}$. This added flexibility enables the modeling of subgroups based on protected attributes that are not explicitly observed but can be inferred through covariates and labels (Romano et al., 2020a). We name our method Kandinsky Conformal Prediction [^1] We note that the name Kandinsky Conformal Prediction coincides with that of an unrelated conformal prediction method for image segmentation (Brunekreef et al., 2024), which balances pixel-level and image-level calibration. Our technique instead focuses on conditional guarantees of general conformal prediction., drawing inspiration from Wassily Kandinsky’s compositions of overlapping geometric forms (see Figure 2), as we provide conditional coverage guarantees for groups that are flexible, overlapping, and probabilistic.
@@
To complement our main algorithm, which ensures training-conditional validity, we also provide a test-time inference algorithm that obtains expected marginal coverage that also takes randomness over the draws on calibration data. By extending Gibbs et al. (2023)’s approach to accommodate group functions defined over both covariates and labels, we obtain an expected coverage error bound of $\mathcal{O}(n_{{\cal G}}^{-1})$. However, this result comes at the cost of computational efficiency, as the algorithm needs to perform test-time inference that solves an instance of quantile regression for every new test example.
@@
    We evaluate our algorithm on real-world tasks, including income prediction, toxic comment detection, and multiple-choice question answering. Our results show that Kandinsky Conformal Prediction consistently achieves the best-calibrated conditional coverage under diverse model architectures, including boosting trees, neural networks, and LLMs, while scaling effectively with the number of groups.
@@ 1.1 Related Work
We introduce Kandinsky Conformal Prediction as an extension of the Mondrian approach, providing coverage guarantees for overlapping group structures. Prior work on overlapping group-conditional coverage primarily considers groups defined solely by covariates $X_{n+1}$. Barber et al. (2021) addresses this by assigning the most conservative prediction set to points in multiple groups. Our algorithm builds on the quantile regression techniques of Jung et al. (2023) and Gibbs et al. (2023), which compute per-example thresholds on non-conformity scores. Blot et al. (2024) extends Gibbs et al. (2023)’s technique to include group membership functions depending on both the label and the covariate, and they establish a one-sided marginal coverage bound. We further provide two-sided training conditional bounds with a different algorithm. Additionally, Areces et al. (2024) establishes a lower bound on group-conditional coverage error rates. Since we generalize to group functions that depend on both covariates and labels, their bound applies to our setting and shows that our error rate is minimax-optimal.
@@
Finally, our work advances the study of conformal prediction under distribution shifts between calibration and test data. Existing research largely focuses on specific cases, such as covariate shift (Tibshirani et al., 2019; Qiu et al., 2023; Yang et al., 2024; Pournaderi and Xiang, 2024) or label shift (Podkopaev and Ramdas, 2021). Plassier et al. (2024) considers general shifts on both the covariate and label distributions. However, all of these approaches are designed for a single target distribution. By introducing a more general class of group functions that depend jointly on covariates and labels, we establish coverage guarantees that remain valid under broader distribution shifts over $\cal X\times\cal Y$, including multiple potential target distributions.
@@ 2 Preliminaries
[…]
Let $(1-\alpha)$ be the target coverage level, where $\alpha\in(0,1)$. We can now restate the definition of training conditional validity to account for the randomness of the algorithm and the prediction set function. Specifically, this requires that for any $f$ the training algorithm returns a (potentially randomized) prediction set function ${\cal C}$ such that, with probability $1-\delta$ over the draws of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the randomness of the conformal prediction algorithm, captured by $\{\varepsilon_{i}\}_{i\in[n]}$, the following holds:
$\mathbb{P}\left[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid\{(X_{i},Y_{i},\varepsilon_{i})\}_{i\in[n]}\right]\approx 1-\alpha,$ (2)
for some $\delta$ in (0,1)(0,1)( 0 , 1 ).
We use weight functions $w:\cal X\times\cal Y\to\mathbb{R}$ to model our conditional coverage guarantees. For any weight function $w$, we measure the error of a fixed prediction set function ${\cal C}$ by its weighted coverage deviation:
$\operatorname{wCD}({\cal C},\alpha,w)\coloneqq\mathbb{E}[w(X,Y)(\mathbbm{1}\{Y\in{\cal C}(X;\varepsilon)\}-(1-\alpha))],$
where the expectation is taken over the test point $(X,Y)$, drawn from ${\cal D}$, and $\varepsilon$, the internal randomness of ${\cal C}$.
@@
Let $\alpha$ be a parameter in (0,1)(0,1)( 0 , 1 ), and let $\{s_{i}\}_{i\in[n]}$ be a set of $n$ points, where for all $i\in[n]$, $s_{i}\in\mathbb{R}$. Then, the largest $\theta^{*}\in\mathbb{R}$ such that
@@ 3 Kandinsky Conformal Prediction
Given the predictor $f$, we use a non-conformity score function $S:{\cal X}\times{\cal Y}\to\mathbb{R}$ to measure how close the prediction $f(x)$ is to the label $y$ for any data point $(x,y)\in{\cal X}\times{\cal Y}$. As is common in conformal prediction, our method ensures the desirable coverage for any score function $S$. However, in our more general framework, we allow for the use of a randomized score function to break potential ties in quantile regression, while keeping the assumptions about distribution ${\cal D}$ minimal. We implement this through a randomized score function ${\widetilde{S}}:{\cal X}\times{\cal Y}\times\mathbb{R}\to\mathbb{R}$ that takes as input the covariates, labels and some random noise $\varepsilon\in\mathbb{R}$.
Our method consists of two steps: one is performed during training, and the other at test time. Similar to Jung et al. (2023), the first step is a quantile regression on the scores of the $n$ data points in the calibration dataset, $\{S(X_{i},Y_{i})\}_{i\in[n]}$. For randomized score functions, we work with the set of randomized scores $\{{\widetilde{S}}(X_{i},Y_{i},\varepsilon_{i})\}_{i\in[n]}$, where for every point $i$, $\varepsilon_{i}$ is drawn independently from a distribution ${\cal D}_{\text{rn}}$. In Algorithm 1, we compute a $(1-\alpha)$-“quantile function” $\widehat{q}$. This $\widehat{q}$ is a weight function from ${\cal W}$ that minimizes the average pinball loss of the $n$ (randomized) scores with parameter $\alpha$.
@@
${\cal C}(X_{n+1};\varepsilon_{n+1})=\left\{y:{\widetilde{S}}(X_{n+1},y,\varepsilon_{n+1})\leq\widehat{q}(X_{n+1},y)\right\}.$
Return: ${\cal C}(X_{n+1};\varepsilon_{n+1})$
Our main result is stated in Theorem 3.1, where we prove that, under the assumption that the distribution of the randomized score ${\widetilde{S}}$ conditioned on the value of the basis $\Phi$ is continuous, the weighted coverage deviation of ${\cal C}$ converges to zero with high probability at a rate of $\mathcal{O}(\sqrt{d/n}+d/n)$, where $d$ is the dimension of $\Phi$. When the distribution of $S(X,Y)\mid\Phi(X,Y)$ is continuous, this theorem holds for the deterministic score function $S$, since setting ${\widetilde{S}}(x,y,\varepsilon)=S(x,y)$ satisfies the assumptions.
@@
Let parameters $\alpha,\delta\in(0,1)$, and ${\cal W}=\{\Phi(\cdot)^{T}\beta:\beta\in\mathbb{R}^{d}\}$ denote a class of linear weight functions over a bounded basis $\Phi:{\cal X}\times{\cal Y}\to\mathbb{R}^{d}$. Without loss of generality, assume $\|\Phi(x,y)\|_{\infty}\leq 1$ for all $x,y$. Assume that the data $\{(X_{i},Y_{i})\}_{i\in[n]}$ are drawn i.i.d. from a distribution ${\cal D}$, $\{\varepsilon_{i}\}_{i\in[n]}$ are drawn i.i.d. from a distribution ${\cal D}_{\text{rn}}$, independently from the dataset, and the distribution of ${\widetilde{S}}\left(X,Y,\varepsilon\right)\mid\Phi(X,Y)$ is continuous. There exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$ given by Algorithms 1 and 2 satisfies, for every $w_{\beta}=\Phi(\cdot)^{T}\beta$,
$\left|\operatorname{wCD}({\cal C},\alpha,w_{\beta})\right|\leq\|\beta\|_{1}\left(C\sqrt{\frac{d}{n}}+\frac{d}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4d/\delta)}{n}}\right).$
As a sketch of the proof, we first establish a connection between the subgradient of the empirical pinball loss and the empirical weighted coverage. We then prove a concentration result for the weighted coverage. In contrast, Roth (2022); Jung et al. (2023) derive concentration results directly for the pinball loss, which leads to a slower convergence rate of $\mathcal{O}((d\log n/n)^{1/4})$.
@@
Let $\alpha,\delta,{\cal W},w_{\beta}$ be specified as in Theorem 3.1 and consider the same assumptions on $\{(X_{i},Y_{i},\varepsilon_{i})\}_{i\in[n]}$, $\widetilde{S}$, and $\Phi$ in Theorem 3.1. Then, there exists an absolute constant $C$ such that the (randomized) prediction set ${\cal C}$, given by Algorithms 1 and 2, satisfies
$\mathbb{E}_{D,E}\left[\sup_{w_{\beta}\in{\cal W}}\frac{\operatorname{wCD}({\cal C},\alpha,w_{\beta})}{\|\beta\|_{1}}\right]\leq C\sqrt{\frac{d}{n}}+\frac{d}{n},$
@@ 3.1 Applications
Suppose we want to achieve the target coverage $(1-\alpha)$ for every group of points $G\subset{\cal X}\times{\cal Y}$ within a finite set of potentially overlapping groups ${\cal G}$. Let $\pi_{G}({\cal C})$ be the probability that the true label is included in a fixed prediction set ${\cal C}$, conditioned on the datapoint $(X_{n+1},Y_{n+1})$ belonging to some group $G$. Formally, with the calibration dataset $D=\{(X_{i},Y_{i})\}_{i\in[n]}$ and the corresponding noise $E=\{\varepsilon_{i}\}_{i\in[n]}$, let
$\pi_{G}({\cal C})\coloneqq\mathbb{P}[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid D,E,(X_{n+1},Y_{n+1})\in G].$
We want to ensure that for every $G\in{\cal G}$, $\pi_{G}({\cal C})\approx 1-\alpha$. Here we can define $\Phi$ to be a $|{\cal G}|$-dimensional vector that has an entry $\mathbbm{1}\{(x,y)\in G\}$ for every $G\in{\cal G}$. Corollary 3.3 provides the rate at which $\pi_{G}({\cal C})$ converges to $(1-\alpha)$, when ${\cal C}$ is constructed by Algorithms 1 and 2. For conciseness, we define $\pi_{G}\coloneqq\mathbb{P}[(X_{n+1},Y_{n+1})\in G]$.
@@
Let parameters $\alpha,\delta\in(0,1)$, and ${\cal W}=\{\sum_{G\in{\cal G}}\beta_{G}\mathbbm{1}\{(x,y)\in G\}:\beta_{G}\in\mathbb{R},\forall G\in{\cal G}\}$. Assume $\{(X_{i},Y_{i},\varepsilon_{i})\}_{i\in[n+1]}$ are i.i.d. samples. Under the same assumptions on $\widetilde{S}$ and $\Phi$ in Theorem 3.1, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$ given by Algorithms 1 and 2 satisfies, for all $G\in{\cal G}$,
$\left|\pi_{G}({\cal C})-(1-\alpha)\right|\leq{}\frac{1}{\pi_{G}}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4|{\cal G}|/\delta)}{n}}\right).$
@@
A special case of group-conditional conformal prediction is Mondrian conformal prediction, where groups in ${\cal G}$ are disjoint and form a partition of the domain ${\cal X}\times{\cal Y}$. Since each $(X_{i},Y_{i})$ in the calibration dataset belongs to exactly one group in ${\cal G}$, we can verify that Algorithm 1 computes a value $\beta_{G}\in\mathbb{R}$ for every group $G$ independently. By Lemma 2.3, for every $G\in{\cal G}$, if $\beta_{G}$ is the largest value that minimized the pinball loss, then it is a $(1-\alpha)$-quantile of the scores of the calibration datapoints that belong to $G$. As a result, for a test point $X_{n+1}$, we construct the prediction set ${\cal C}(X_{n+1})$ by including every $y$ where the score ${\widetilde{S}}(X_{n+1},y,\varepsilon)$ is at most the quantile $\beta_{G}$ corresponding to the group $G$ that contains $(X_{n+1},y)$. This method is the same as the one presented in Vovk et al. (2003), with the difference that in that method, the quantiles are slightly adjusted to achieve the expected marginal coverage guarantee through exchangeability.
@@
$\pi^{Z}_{G}({\cal C})\coloneqq\mathbb{P}[Y_{n+1}\in{\cal C}(X_{n+1};\varepsilon_{n+1})\mid D,E,Z_{n+1}\in G].$
Then, we want to ensure that $\pi^{Z}_{G}({\cal C})\approx 1-\alpha$ for every $G\in{\cal G}$. To achieve this, we set the basis of the weight function class as the probability of $Z\in G$ given a statistic $\phi(X,Y)$. For every $G\in{\cal G}$,
@@
For our theoretical result, we assume that we already know these conditional probabilities, $\Phi_{G}(x,y)$, for all $G\in{\cal G}$, and that the calibration and the test data do not include unobserved attributes. Corollary 3.4 shows that by constructing ${\cal C}$ according to Algorithms 1 and 2, the coverage converges to $(1-\alpha)$ at an $O(\sqrt{|{\cal G}|/n}+|{\cal G}|/n)$ rate. For brevity, we define $\pi_{G}^{Z}\coloneqq\mathbb{P}[Z_{n+1}\in G]$.
@@
Let $\alpha,\delta\in(0,1)$, $\phi(X,Y)$ be a sufficient statistic for ${\widetilde{S}}$, such that ${\widetilde{S}}$ is conditionally independent of $X,Y$ given $\phi$, and ${\cal W}=\{\sum_{G\in{\cal G}}\beta_{G}\Phi_{G}:\beta_{G}\in\mathbb{R},\forall~G\in{\cal G}\}.$ Assume $\{(X_{i},Y_{i},\varepsilon_{i})\}_{i\in[n+1]}$ are i.i.d. samples. Under the same assumptions on $\widetilde{S}$ and $\Phi$ in Theorem 3.1, there exists an absolute constant $C$ such that, with probability at least $1-\delta$ over the randomness of the calibration dataset $\{(X_{i},Y_{i})\}_{i\in[n]}$ and the noise $\{\varepsilon_{i}\}_{i\in[n]}$, the (randomized) prediction set ${\cal C}$ given by Algorithms 1 and 2 satisfies, for all $G\in{\cal G}$,
$\left|\pi^{Z}_{G}({\cal C})-(1-\alpha)\right|\leq\frac{1}{\pi_{G}^{Z}}\left(C\sqrt{\frac{|{\cal G}|}{n}}+\frac{|{\cal G}|}{n}+\max\{\alpha,1-\alpha\}\sqrt{\frac{2\ln(4|{\cal G}|/\delta)}{n}}\right).$